OpenStudy (anonymous):

find the limit x^(-17/lnx) x->0+

OpenStudy (anonymous):

Try taking the log

OpenStudy (anonymous):

\[x ^{-17/lnx} x->0^{+}\]

OpenStudy (anonymous):

taking the log of lnx, or the whole equation? I don't know if this makes a difference, but were using Hopitals rule

OpenStudy (anonymous):

you can just take the log of the whole equation. ln(x^(-17/ln(x))=-17.

OpenStudy (anonymous):

thus the equation is equal to the constant e^(-17).

OpenStudy (anonymous):

thus the limit is also e^(-17)

OpenStudy (anonymous):

the derivative of lnx is (1/x) how were you able to go from this equation to base e. It shows the same steps in my book, but I am just a little confused to how you got there

OpenStudy (anonymous):

(thank you so much for your help by the way)

OpenStudy (anonymous):

1) ln(a^b)=b*ln(a) 2) ln(e)=1 3) e^ln(x)=x

OpenStudy (anonymous):

so if ln(x^(-17/lnx))=-17, then x^(-17/lnx)=e^(-17)

OpenStudy (anonymous):

you literally explained it so simply that my professor can't even sum it up that way

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

i really appreciate this explanation, thank you so much!